The equation y = 5Sin (3x -4t), where y is in millimetres, x is in metres and t is in seconds represents a wave motion. Determine the frequency of the wave



1
Expert's answer
2022-02-18T08:36:30-0500

Solution

Compare to standard equation

"y=A\sin(kx-wt)"

So frequency

"f=\frac{w}{2\pi}\\=\frac{4}{2\times3.14}\\=0.64Hz"